Grade 12Mechanics

a particle is moving in a straight line whose position x in time t is described by x= t³-t² . then the average acceleration from t= 2 to t=4 is

janhvi singh
9 Years agoGrade 12

1 Answer

Aman gohel
9 Years ago
X=t^3-t^2. =>dX/dt=d(t^3-t^2)/dt =>v=3t^2-2t =>dV/dt=d(3t^2-2t)/dt =>a=6t-2 At time t=2s, a=6×2-2=10m/s^2. At time t=4s a=4×6-2=22m/s^2change in a=22-10=12m/s^2 Therefore average acceleration=12/2=6m/s^3(Ans)